OpenStudy (anonymous):

How to solve for 3 over 4 * x + 4 = 10 ?

OpenStudy (anonymous):

3/4(x+4)=10 is this the problem?

OpenStudy (e.cociuba):

1.) Subtract 4 from 10

OpenStudy (e.cociuba):

You will get wat??

OpenStudy (anonymous):

-6

OpenStudy (e.cociuba):

Take -6 and divide it by 3/4x and wat do u get.....

OpenStudy (e.cociuba):

You get x= -4.5

OpenStudy (anonymous):

3/4x + 4 = 10 subtract 4 from both sides 3/4x = 6